# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ 1.2.1 Flexible nonstationary max-id spatial extremes model
Our model builds upon the max-id process introduced by Reich and Shaby (2012) and extended by Bopp et al. (2021). Importantly, a novel extension of our model is its ability to realistically capture both short-range asymptotic dependence (AD), mid-range asymptotic independence (AI), and long-range exact independence, as explained in more detail in Section 1.3, and it can accommodate nonstationary in space and time. As in these earlier works, we start by defining the spatial observation model as
@@
where $\mathcal{S}\in\mathbb{R}^{2}$ is the domain of interest and $\epsilon(\boldsymbol{s})$ is a noise process with independent Fréchet$(0,\tau,1/\alpha_{0})$ marginal distributions; that is, $\Pr\{\epsilon(\boldsymbol{s})\leq x\}=\exp\{-(x/\tau)^{1/\alpha_{0}}\}$, where $x>0$, $\tau>0$ and $\alpha_{0}>0$. Then, $Y(\boldsymbol{s})$ is constructed using a low-rank representation:
$Y(\boldsymbol{s})=\left\{\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s})^{\frac{1}{\alpha}}Z_{k}\right\}^{\alpha_{0}},$ (4)
@@
where $\alpha\in(0,1)$, $\{\omega_{k}(\boldsymbol{s}):k=1,\ldots,K\}$ are fixed compactly-supported radial basis functions centered at $K$ pre-specified knots such that $\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s})=1$ for any $\boldsymbol{s}\in\mathcal{S}$, and $\{Z_{k}:k=1,\ldots,K\}$ are latent variables independently distributed as exponentially-tilted positive-stable (PS) random variables (Hougaard, 1986). We write $Z_{k}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\mathrm{expPS}(\alpha,\gamma_{k}),\;k=1,\ldots,K$, in which $\alpha\in(0,1)$ determines the rate at which the power-law tail of $\mathrm{expPS}(\alpha,0)$ tapers off, and the tilting parameters $\gamma_{k}\geq 0$ determine the extent of tilting, with larger values of $\gamma_{k}$ leading to lighter-tailed $Z_{k}$ (see Section B.1 of the Supplementary Material).
Our model, inspired by Reich and Shaby (2012) and Bopp et al. (2021), presents a novel class of spatial extremes models. In Reich and Shaby (2012) and Bopp et al. (2021), the basis functions lack compact support and all tilting parameters are fixed at either $\gamma_{k}\equiv 0$ or $\gamma_{k}\equiv\gamma>0$, resulting in only AI or AD for all pairs of locations, respectively. In contrast, the use of compactly-supported basis function and spatially-varying tilting parameters creates a spatial-scale aware extremal dependence model, enabling local asymptotic dependence (AD) or asymptotic independence (AI) for nearby locations while ensuring long-range AI for distant locations, which is a significant advancement in spatial extremes.
Moreover, while both previous works use a noise process with Fréchet$(0,1,1/\alpha)$ marginals (i.e., $\alpha_{0}=\alpha$), our approach decouples noise variance from tail heaviness, providing better noise control for each time point, while keep the appealing property of being max-id as shown below in Section 1.3. Additionally, we embed our model in a VAE, leveraging variational-Bayes-based parameter estimation for rapid computation, modeling, and emulation of high-dimensional spatial extremes (on the order of more than 10,000 locations) without distributing the model parameter estimation over a number of subdomains. Furthermore, we allow different concentration parameters $\alpha_{t}$ and tilting parameters $\boldsymbol{\gamma}_{t}=\{\gamma_{kt}:k=1,\ldots,K\}$, at different time points $t=1,\ldots,n_{t}$, which automatically ensures a temporally non-stationary spatial extremes model.
To the best of our knowledge, our XVAE is the first attempt to capture both spatially and temporally varying extremal dependence structures simultaneously in one model at the scale that we consider here. Zhong et al. (2022) achieved it at a much smaller scale and using a quite rigid covariate-based approach. In Section A of the Supplementary Material, we compare our max-id model with other existing models more in depth.
1.2.2 XVAE: A VAE incorporating the proposed max-id spatial model
@@
For notational simplicity hereafter, we denote by $\boldsymbol{X}_{t}=\{X_{t}(\boldsymbol{s}_{j}):j=1,\ldots,n_{s}\}$ the observations of process (3) at time $t=1,\ldots,n_{t}$, and by $\boldsymbol{Z}_{t}=\{Z_{kt}:k=1,\ldots,K\}$ the corresponding latent variables.
@@
Inference for our flexible extremes model on large spatial datasets poses challenges. A streamlined Metropolis–Hastings MCMC algorithm would be time-consuming and hard to monitor when confronted with the scale of our spatial data in Section 4, where a considerable number of local basis functions $K$ is necessary to capture intricate local extremes accurately. Additionally, when there are many time replicates, inferring the parameters $(\alpha_{t},\boldsymbol{\gamma}_{t},\boldsymbol{Z}_{t})$ for all time points becomes extremely challenging. Moreover, with a large sample size $n_{s}$, evaluating the likelihood $p_{\boldsymbol{\theta}}(\boldsymbol{x}\mid\boldsymbol{z})$ for a single iteration becomes computationally expensive, and updating different parameters iteratively further exacerbates this issue. To overcome these challenges, we employ the encoding-decoding VAE paradigm described in Section 1.1 and modify it to account for our extremes framework. For $t=1,\ldots,n_{t}$, our encoder $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ maps each observed replicate $\boldsymbol{x}_{t}$ to the latent space and allows fast random sampling of $\{\boldsymbol{Z}_{t}^{1},\ldots,\boldsymbol{Z}_{t}^{L}\}$ that will be approximately distributed according to the true posterior $p_{\boldsymbol{\theta}_{t}}(\cdot\mid\boldsymbol{x}_{t})$ because of the ELBO regularization, in which $\boldsymbol{\theta}_{t}=(\alpha_{0},\tau,\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}$; recall Eqs. (1) and (2). The details of this approach are provided below.
@@
$\begin{split}\boldsymbol{z}_{t}&=\boldsymbol{\mu}_{t}+\boldsymbol{\zeta}_{t}\odot\boldsymbol{\eta}_{t},\\ \eta_{kt}&\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Normal}(0,1),\\ (\boldsymbol{\mu}_{t}^{T},\log\boldsymbol{\zeta}_{t}^{T})^{T}&=\mathrm{EncoderNeuralNet}_{\boldsymbol{\phi}_{e}}(\boldsymbol{x}_{t}),\end{split}$ (5)
where we use a standard reparameterization trick with an auxiliary variable $\boldsymbol{\eta}_{t}=\{\eta_{kt}:k=1,\ldots,K\}$ and $\odot$ is the elementwise product (Kingma and Welling, 2013). This trick enables fast computation of Monte Carlo estimates of $\nabla_{\boldsymbol{\phi}_{e}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}$, the gradient of the ELBO with respect to $\boldsymbol{\phi}_{e}$ (see Section D.1 for details). Also, by controlling the mean $\boldsymbol{\mu}_{t}$ and variance $\boldsymbol{\zeta}^{2}_{t}$, the distributions $q_{\phi_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ are enforced to be close to $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ for each $t$. This is the primary role of the deep neural network in (5) — i.e., to learn the complex relationship between the inputs $\boldsymbol{x}_{t}$ and the latent process, $\boldsymbol{z}_{t}$. The specific neural network architecture and implementation details are given in Section D of the Supplementary Material.
@@
This is determined by our model construction. Specifically, denoting the density function of the $\mathrm{expPS}(\alpha,\gamma_{k})$ distribution by $h(z;\alpha,\gamma_{k})$, the prior on $\boldsymbol{z}_{t}$ can be written as
@@
which depends on parameters $(\alpha_{t},\gamma_{kt}:t=1,\ldots,T;k=1,\ldots,K)$.
@@
Our decoder is based on the flexible max-id spatial extremes model described in Section 1.2.1. Specifically, recall from Eq. (4) that $\Pr(\boldsymbol{X}_{t}\leq\boldsymbol{x}_{t}\mid\boldsymbol{Z}_{t}=\boldsymbol{z}_{t})=\exp\{-\sum_{j=1}^{n_{s}}\left({\tau}/{x_{jt}}\right)^{{1}/{\alpha_{0}}}\sum_{k=1}^{K}\omega_{kj}^{{1}/{{\alpha}_{t}}}z_{kt}\}$. Differentiating this conditional distribution function gives the exact form of the decoder:
$p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t}\mid\boldsymbol{z}_{t})=\left(\frac{1}{\alpha_{0}}\right)^{n_{s}}\left\{\prod_{j=1}^{n_{s}}\frac{1}{x_{jt}}\left(\frac{x_{jt}}{\tau y_{jt}}\right)^{-1/\alpha_{0}}\right\}\exp\left\{-\sum_{j=1}^{n_{s}}\left(\frac{x_{jt}}{\tau y_{jt}}\right)^{-1/\alpha_{0}}\right\},$ (7)
where $y_{jt}=\sum_{k=1}^{K}\omega_{kj}^{1/{\alpha}_{t}}z_{kt}$. This distribution depends on the Fréchet parameters $(\alpha_{0},\tau)^{T}$ and the dependence parameters $(\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}$ inherited from the prior distribution of $\boldsymbol{z}_{t}$. The decoder neural network estimates these dependence parameters as
$(\hat{\alpha}_{t},\hat{\boldsymbol{\gamma}}_{t}^{T})^{T}=\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d,0}}(\boldsymbol{Z}_{t})$ (8)
where $\boldsymbol{\phi}_{d,0}$ are the bias and weight parameters of this neural network (see Eqs. (D.1) and (D.3) of the Supplementary Material for more details). Combining $\boldsymbol{\phi}_{d,0}$ with the Fréchet parameters $(\alpha_{0},\tau)^{T}$, we write $\boldsymbol{\phi}_{d}=(\alpha_{0},\tau,\boldsymbol{\phi}_{d,0}^{T})^{T}$. We must use the variational procedure to find estimates of parameters $\boldsymbol{\phi}_{d}$ and the encoder neural network parameters, $\boldsymbol{\phi}_{e}$.
@@
By drawing $L$ independent samples ${\boldsymbol{Z}_{t}^{1},\ldots,\boldsymbol{Z}_{t}^{L}}$ using Eq. (5), we can derive the Monte Carlo estimate of the ELBO, $\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$, as in Eq. (2). In Section D.3 of the Supplementary Material, we detail the stochastic gradient search used to find the $\boldsymbol{\phi}_{e}$ and $\boldsymbol{\phi}_{d}$ that maximize $\sum_{t=1}^{n_{t}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$. We stress that our XVAE is an example of “amortized inference” (Zammit-Mangion et al., 2024): there is a substantial training cost up front, but once the XVAE is trained, posterior simulation of new latent variables $\boldsymbol{Z}_{t}$ can be performed very efficiently following Eq. (5) and synthetic data can be generated extremely quickly by passing them through the decoder (8) and sampling from the model $p_{\hat{\boldsymbol{\theta}}_{t}}(\boldsymbol{x}\mid\boldsymbol{Z}_{t})$ specified by Eqs. (3) and (4), in which $\hat{\boldsymbol{\theta}}_{t}=(\hat{\alpha}_{0},\hat{\tau},\hat{\alpha}_{t},\hat{\boldsymbol{\gamma}}_{t}^{T})^{T}$.

### v4

@@ 3.1 Flexible nonstationary max-id spatial extremes model
Our model builds upon the max-id process proposed by Reich and Shaby (2012) and extended by Bopp et al. (2021). Importantly, a novel extension of our model is its ability to realistically capture the change of asymptotic dependence class as a function of distance, as explained in more detail in Section 3.1.1, and to accommodate nonstationarity in space and time. Similar to these earlier works, we define the spatial observation model as
@@
where $\mathcal{S}\in\mathbb{R}^{2}$ is the domain of interest and $\epsilon(\boldsymbol{s})$ is a noise process with independent Fréchet$(0,\tau,1/\alpha_{0})$ marginal distributions; that is, $\Pr\{\epsilon(\boldsymbol{s})\leq x\}={\color[rgb]{0,0,0}\exp\{-(x/\tau)^{-1/\alpha_{0}}\}}$, where $x>0$, $\tau>0$ and $\alpha_{0}>0$. Then, $Y(\boldsymbol{s})$ is constructed using a low-rank representation:
$Y(\boldsymbol{s})=\left\{\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s})^{1/\alpha}Z_{k}\right\}^{\alpha_{0}},$ (4)
@@
where $\alpha\in(0,1)$, $\{\omega_{k}(\boldsymbol{s}):k=1,\ldots,K\}$ are fixed compactly-supported radial basis functions centered at $K$ pre-specified knots such that $\sum_{k=1}^{K}\omega_{k}(\boldsymbol{s})=1$ for any $\boldsymbol{s}\in\mathcal{S}$, and $\{Z_{k}:k=1,\ldots,K\}$ are independently distributed as exponentially-tilted positive-stable (expPS) random variables, whose densities are of the form
$h(x;\alpha,\gamma_{k})=\frac{f_{\alpha}(x)\exp(-\gamma_{k}x)}{\exp(-\gamma_{k}^{\alpha})},\;x>0,\;k=1,\ldots,K;$ (5)
here, $f_{\alpha}$ is the density function of a positive-stable variable defined through its Laplace transform $\int_{\mathbb{R}}\exp(-sx)f_{\alpha}(x){\rm d}x=\exp(-s^{\alpha}),s\geq 0$ (Hougaard, 1986), $\alpha\in(0,1)$ determines the rate at which the power-law tail of $f_{\alpha}$ tapers off, and the tilting parameters $\gamma_{k}\geq 0$ determine the extent of tilting, with larger values of $\gamma_{k}$ leading to lighter-tailed $Z_{k}$; see Section A.1 of the Supplementary Material for details. We write $Z_{k}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\mathrm{expPS}(\alpha,\gamma_{k})$.
Our spatial extremes model, while inspired from Reich and Shaby (2012) and Bopp et al. (2021), has several key novelties. In both Reich and Shaby (2012) and Bopp et al. (2021), the basis functions lack compact support and all tilting parameters are fixed at either $\gamma_{k}\equiv 0$ or $\gamma_{k}\equiv\gamma>0$, resulting in only AD or AI for all pairs of locations, respectively. By contrast, in our model, the use of compactly-supported basis functions and spatially-varying tilting parameters creates a spatial-scale aware extremal dependence model, which enables us to capture local AD or AI for nearby locations while ensuring long-range AI for distant locations—a significant advancement in the spatial extremes literature. Moreover, while both previous works use a noise process with Fréchet$(0,1,1/\alpha)$ marginals (i.e., setting $\alpha_{0}=\alpha$), our approach decouples the noise variance from the tail heaviness, providing better noise control for each time point, while keeping the appealing property of max-infinite divisibility as shown in Section 3.1.1. Finally, when temporal replicates are available, we shall allow the concentration parameter $\alpha$ and tilting parameters $\boldsymbol{\gamma}=\{\gamma_{k}:k=1,\ldots,K\}$ to take different values across time points (i.e., allowing such parameters, denoted by $\alpha_{t}$ and $\boldsymbol{\gamma}_{t}=\{\gamma_{kt}:k=1,\ldots,K\}$, respectively, to change over time $t$), thus making our spatial extremes model nonstationary over both space and time. To the best of our knowledge, this is the first attempt to capture both spatially and temporally varying extremal dependence structures simultaneously in one model, without sub-domain partitioning, at the spatio-temporal scale that we consider here. Zhong et al. (2022) achieved it at a much smaller scale and using quite a rigid covariate-based approach to capture nonstationarity.
3.1.1 Marginal and dependence properties
In this section, we examine the marginal and joint tail behavior of the spatial model (3). When temporal replicates are available, one can readily replace the parameters $\alpha$ and $\gamma_{k}$ with temporally-varying parameters, $\alpha_{t}$ and $\gamma_{kt}$, respectively. For notational simplicity, we write $X_{j}=X(\boldsymbol{s}_{j})$, $\omega_{kj}=\omega_{k}(\boldsymbol{s}_{j})$, $k=1,\ldots,K$, $j=1,\ldots,n_{s}$, with $n_{s}$ the number of observed locations, and define $\mathcal{C}_{j}=\{k:\omega_{kj}\neq 0,k=1,\ldots,K\}$. We require that any location $\boldsymbol{s}\in\mathcal{S}$ be covered by at least one basis function, so $\mathcal{C}_{j}$ cannot be empty for any $j$.
We first study the marginal distributions of the process (3).
Proposition 3.1.
Let $\mathcal{D}=\{k:\gamma_{k}=0,\;k=1,\ldots,K\}$ and $\bar{\mathcal{D}}$ be the complement of $\mathcal{D}$. For the process (3), the marginal distribution function of $X_{j}=X(\boldsymbol{s}_{j})$ can be written as
$F_{j}(x)=\exp\left\{\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left(\gamma_{k}+{\tau}^{{1/\alpha_{0}}}\omega_{kj}^{{1/\alpha}}x^{-{1/\alpha_{0}}}\right)^{\alpha}\right\}.$ (6)
As $x\rightarrow\infty$, the survival function $\bar{F}_{j}(x)=1-F_{j}(x)\sim c_{j}(x/\tau)^{-{1/\alpha_{0}}}$ if $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, and $\bar{F}_{j}(x)\sim c_{j}^{\prime}(x/\tau)^{-{\alpha/\alpha_{0}}}$ if $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, where $c_{j}=\alpha\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha-1}\omega_{kj}^{1/\alpha}$, $c^{\prime}_{j}=\sum_{k\in\mathcal{D}}\omega_{kj}$.
The proof of this result can be found in Section A.2 of the Supplementary Material. It indicates that the process (3) has Pareto-like marginal tails at any location in the domain $\mathcal{S}$. If $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, that is, if the $j$th location is impacted by an “un-tilted knot” (i.e., a knot with $\gamma_{k}=0$ in the $\mathrm{expPS}(\alpha,\gamma_{k})$ distribution of the corresponding latent variable $Z_{k}$), then $\bar{F}_{j}(x)=O(x^{-{\alpha/\alpha_{0}}})$ as $x\rightarrow\infty$ since $\alpha\in(0,1)$. If, however, the location is not within the reach of an un-tilted knot, then the marginal distribution is less heavy-tailed.
To derive the extremal dependence structure, we first calculate the joint distribution function of a $n_{s}$-variate random vector $(X_{1},\ldots,X_{n_{s}})^{\rm T}$ drawn from the process (3).
Proposition 3.2.
Under the definitions and notation as established in Proposition 3.1, for locations $\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n_{s}}\in\mathcal{S}$, the exact form of the joint distribution function of the random vector $(X_{1},\ldots,X_{n_{s}})^{\rm T}$ can be written as
$F(x_{1},\ldots,x_{n_{s}})=\exp\left\{\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left(\gamma_{k}+\tau^{{1/\alpha_{0}}}\sum_{j=1}^{n_{s}}\omega_{kj}^{{1/\alpha}}x_{j}^{-{1/\alpha_{0}}}\right)^{\alpha}\right\}.$ (7)
The proof of Proposition 3.2 is given in Section A.3 of the Supplementary Material. Eq. (7) ensures that $F^{1/r}(x_{1},\ldots,x_{n_{s}})$ is a valid distribution function on $\mathbb{R}^{n_{s}}$ for any real $r>0$, of the same form as (7) but with tilting indices $\{\gamma_{1}r^{-1/\alpha},\ldots,\gamma_{K}r^{-1/\alpha}\}$ and scale parameter $\tau\,r^{-\alpha_{0}/\alpha}$. By definition, the process $\{X_{t}(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\}$ is thus max-infinitely divisible (max-id). It becomes max-stable only when it remains within the same location-scale family, i.e., when $\gamma_{1}=\cdots=\gamma_{K}=0$.
We now characterize the tail dependence structure of model (3) using both $\chi_{ij}$ defined in Eq. (1) and the complementary measure $\eta_{ij}$ defined by $\Pr\{X_{i}>F_{i}^{-1}(u),X_{j}>F_{j}^{-1}(u)\}=\mathcal{L}\{(1-u)^{-1}\}(1-u)^{1/\eta_{ij}}$, where $\mathcal{L}$ is slowly varying at infinity, i.e., $\mathcal{L}(tx)/\mathcal{L}(t)\rightarrow 1$ as $t\rightarrow\infty$ for all $x>0$. The value of $\eta_{ij}\in(0,1]$ is used to differentiate between the different levels of dependence exhibited by a pair $(X_{i},X_{j})^{\rm T}$. When $\eta_{ij}=1$ and $\mathcal{L}(t)\not\rightarrow 0$ as $t\rightarrow\infty$, $(X_{i},X_{j})^{\rm T}$ is AD ($\chi_{ij}>0$), and the remaining cases are all AI ($\chi_{ij}=0$; see Ledford and Tawn, 1996), with stronger tail dependence for larger values of $\eta_{ij}$.
Theorem 3.3.
Under the assumptions of Propositions 3.1 and 3.2, the process $\{X(\boldsymbol{s})\}$ defined in (3) has a tail dependence structure characterized as follows:
  1. (a)
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, we have $\chi_{ij}=0$ with $\eta_{ij}=1/2$.
  2. (b)
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have $\chi_{ij}=0$ with $\eta_{ij}=\frac{\alpha}{\alpha+1}$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset$ and $\eta_{ij}=1/2$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$.
  3. (c)
    If $\mathcal{C}_{i}\cap\mathcal{D}\neq\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have $\chi_{ij}=2-d_{ij}$ with $\eta_{ij}=1$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, where $d_{ij}=\sum_{k\in\mathcal{D}}\{(\omega_{ki}/c^{\prime}_{i})^{{1/\alpha}}+(\omega_{kj}/c^{\prime}_{j})^{{1/\alpha}}\}^{\alpha}\in(1,2)$, and $\chi_{ij}=0$ with $\eta_{ij}=1/2$ when $\mathcal{C}_{i}\cap\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$.
The proof of this result is given in Section A.4 of the Supplementary Material. The local dependence strength is proportional to the tail-heaviness of the latent variable of the closest knot. There is local AD if $\gamma_{k}=0$, and local AI if $\gamma_{k}>0$, as expected. The sets $\mathcal{C}_{j}\cap\mathcal{D}$, $j=1,\ldots,n_{s}$, are crucial to the behavior of the so-called exponent function in the limiting distribution for normalized maxima. This causes both the asymptotic and sub-asymptotic dependence strength to rely on the tail-heaviness of the local expPS variables and the basis function weights; see Remark 5 in Section A.4 of the Supplementary Material for specifics.
The compactness of the basis functions’ support yields long-range exact independence (thus, also AI) for two far-apart sites that are impacted by disjoint sets of basis functions; this is similar in spirit to the Cauchy convolution process of Krupskii and Huser (2022), though their model construction is different and less computationally tractable than ours.
3.2 XVAE: A VAE incorporating the proposed max-id model
@@
Hereafter, we denote by $\boldsymbol{X}_{t}=\{X_{t}(\boldsymbol{s}_{j}):j=1,\ldots,n_{s}\}$ the realizations of process (3) at time $t=1,\ldots,n_{t}$, and by $\boldsymbol{Z}_{t}=\{Z_{kt}:k=1,\ldots,K\}$ the corresponding latent variables.
@@
Inference for our flexible extremes model on large spatial datasets poses challenges. A streamlined Metropolis–Hastings MCMC algorithm would be time-consuming and hard to monitor when confronted with the scale of our spatial data in Section 5, where a considerable number of local basis functions $K$ is necessary to capture local extremes. Additionally, when there are many time replicates, inferring time-varying parameters $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\top})^{\top}$ and latent variables $\boldsymbol{Z}_{t}$ at all time points becomes extremely challenging. To overcome these challenges, we modify the encoding-decoding VAE paradigm described in Section 2.2 to account for our extremes framework. For $t=1,\ldots,n_{t}$, our encoder $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ maps each observed replicate $\boldsymbol{x}_{t}$ to the latent space and allows fast random sampling of $\{\boldsymbol{Z}_{t}^{1},\ldots,\boldsymbol{Z}_{t}^{L}\}$ that will be approximately distributed according to the true posterior $p_{\boldsymbol{\theta}_{t}}(\cdot\mid\boldsymbol{x}_{t})$ because of the ELBO regularization, in which $\boldsymbol{\theta}_{t}=(\alpha_{0},\tau,\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}$; see Eq. (2). The details of this approach are provided below (see also the illustration in Figure 1).
@@
$\begin{split}\boldsymbol{z}_{t}&=\boldsymbol{\mu}_{t}+\boldsymbol{\zeta}_{t}\odot\boldsymbol{\eta}_{t},\\[-4.30554pt] \eta_{kt}&\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Normal}(0,1),\\[-4.30554pt] (\boldsymbol{\mu}_{t}^{\rm T},\log\boldsymbol{\zeta}_{t}^{\rm T})^{\rm T}&=\mathrm{EncoderNeuralNet}_{\boldsymbol{\phi}_{e}}(\boldsymbol{x}_{t}),\end{split}$ (8)
where $\odot$ is the elementwise product, and a standard reparameterization trick with an auxiliary variable $\boldsymbol{\eta}_{t}=\{\eta_{kt}:k=1,\ldots,K\}$ is used to enable fast computation of Monte Carlo estimates of the gradient $\nabla_{\boldsymbol{\phi}_{e}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}$. Also, by controlling the mean $\boldsymbol{\mu}_{t}$ and variance $\boldsymbol{\zeta}^{2}_{t}$, the distributions $q_{\phi_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ are enforced to be close to $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ for each $t$. This is the primary role of the deep neural network in (8)—to learn the complex relationship between the inputs $\boldsymbol{x}_{t}$ and the latent process $\boldsymbol{z}_{t}$. The specific neural network architecture and implementation details are given in Section D of the Supplementary Material.
@@
This is determined by our model construction. Specifically, the prior on $\boldsymbol{z}_{t}$ can be written as
@@
in which $h(\cdot;\alpha_{t},\gamma_{kt})$ is the density function of $\mathrm{expPS}(\alpha_{t},\gamma_{kt})$, as defined in (5).
@@
Our decoder is based on the flexible max-id spatial extremes model described in Section 3.1. Specifically, recall from Eq. (4) that $\Pr(\boldsymbol{X}_{t}\leq\boldsymbol{x}_{t}\mid\boldsymbol{Z}_{t}=\boldsymbol{z}_{t})=\exp\{-\sum_{j=1}^{n_{s}}\left({\tau}/{x_{jt}}\right)^{{1}/{\alpha_{0}}}\sum_{k=1}^{K}\omega_{kj}^{{1}/{{\alpha}_{t}}}z_{kt}\}$. Differentiating this conditional distribution function gives the exact form of the decoder:
$p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t}\mid\boldsymbol{z}_{t})=\left({1/\alpha_{0}}\right)^{n_{s}}\left\{\prod_{j=1}^{n_{s}}\frac{1}{x_{jt}}\left(\frac{x_{jt}}{\tau y_{jt}}\right)^{-1/\alpha_{0}}\right\}\exp\left\{-\sum_{j=1}^{n_{s}}\left(\frac{x_{jt}}{\tau y_{jt}}\right)^{-1/\alpha_{0}}\right\},$ (10)
where $y_{jt}={\color[rgb]{0,0,0}(\sum_{k=1}^{K}\omega_{kj}^{1/{\alpha}_{t}}z_{kt})^{\alpha_{0}}}$. This distribution depends on the Fréchet parameters $(\alpha_{0},\tau)^{\rm T}$ and the dependence parameters $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}$ inherited from the prior distribution of $\boldsymbol{z}_{t}$. The decoder neural network estimates these dependence parameters as
$(\hat{\alpha}_{t},\hat{\boldsymbol{\gamma}}_{t}^{\rm T})^{\rm T}=\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d,0}}(\boldsymbol{Z}_{t}),$ (11)
where $\boldsymbol{\phi}_{d,0}$ are the bias and weight parameters of this neural network (see Eqs. (D.1) and (D.3) of the Supplementary Material for more details). Combining $\boldsymbol{\phi}_{d,0}$ with the Fréchet parameters $(\alpha_{0},\tau)^{\rm T}$, we write $\boldsymbol{\phi}_{d}=(\alpha_{0},\tau,\boldsymbol{\phi}_{d,0}^{\rm T})^{\rm T}$. We use the variational procedure to find estimates of parameters $\boldsymbol{\phi}_{d}$ and the encoder neural network parameters $\boldsymbol{\phi}_{e}$.
Figure 1: Diagram of a variational autoencoder (VAE) with the reparameterization trick.
@@
By drawing $L$ independent samples ${\boldsymbol{Z}_{t}^{1},\ldots,\boldsymbol{Z}_{t}^{L}}$ using Eq. (8), we can derive the Monte Carlo estimate of the ELBO, $\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$, and then find the parameters $\boldsymbol{\phi}_{e}$ and $\boldsymbol{\phi}_{d}$ that maximize $\sum_{t=1}^{n_{t}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$ via stochastic gradient search, as detailed in Section D of the Supplementary Material. We stress again that our XVAE is a “semi-amortized” inference approach (Zammit-Mangion et al., 2025): there is a substantial training cost up front, but once the XVAE is trained, posterior simulation of new latent variables $\boldsymbol{Z}_{t}$ can be performed very efficiently following Eq. (8) and synthetic data can be generated extremely quickly by passing them through the decoder (11) and sampling from the model $p_{\hat{\boldsymbol{\theta}}_{t}}(\boldsymbol{x}\mid\boldsymbol{Z}_{t})$ specified by Eqs. (3) and (4), in which $\hat{\boldsymbol{\theta}}_{t}=(\hat{\alpha}_{0},\hat{\tau},\hat{\alpha}_{t},\hat{\boldsymbol{\gamma}}_{t}^{\rm T})^{\rm T}$. The XVAE would, however, have to be retrained with new observations $\boldsymbol{X}_{t}$, $t=1,\ldots,n_{t}$.